# DiffOG: Differentiable Policy Trajectory Optimization with Generalizability

### v3

@@ Abstract
Imitation learning-based visuomotor policies excel at manipulation tasks but often produce suboptimal action trajectories compared to model-based methods. Directly mapping camera data to actions via neural networks can result in jerky motions and difficulties in meeting critical constraints, compromising safety and robustness in real-world deployment. For tasks that require high robustness or strict adherence to constraints, ensuring trajectory quality is crucial. However, the lack of interpretability in neural networks makes it challenging to generate constraint-compliant actions in a controlled manner. This paper introduces differentiable policy trajectory optimization with generalizability (DiffOG), a learning-based trajectory optimization framework designed to enhance visuomotor policies. By leveraging the proposed differentiable formulation of trajectory optimization with transformer, DiffOG seamlessly integrates policies with a generalizable optimization layer. DiffOG refines action trajectories to be smoother and more constraint-compliant while maintaining alignment with the original demonstration distribution, thus avoiding degradation in policy performance. We evaluated DiffOG across 11 simulated tasks and 2 real-world tasks. The results demonstrate that DiffOG significantly enhances the trajectory quality of visuomotor policies while having minimal impact on policy performance, outperforming trajectory processing baselines such as greedy constraint clipping and penalty-based trajectory optimization. Furthermore, DiffOG achieves superior performance compared to existing constrained visuomotor policy. Please visit the project website for more details: https://zhengtongxu.github.io/diffog-website/.
@@ II-A Supervised Policy Learning
Supervised policy learning entails training a policy using supervised learning techniques on a dataset of pre-collected demonstrations [15]. Recent advancements in machine learning algorithms and data collection frameworks have substantially advanced the capabilities and effectiveness of supervised policy learning in manipulation applications. For example, various data collection frameworks and devices have been proposed in [6, 35, 8, 21, 20, 24], enabling efficient data collection and the training of effective manipulation policies. To better capture the multi-modal action distributions present in demonstration data, prior works have explored various techniques, such as using diffusion models [5], VQ-VAE [9], and energy-based models [7]. Furthermore, various effective action representations for supervised policy learning have been proposed, such as flow by point tracking [30, 27] and equivariant action representation [25]. Moreover, many of works have contributed to training more general and scalable policy policies on large-scale datasets [13, 14, 3, 11].
@@ III Method
III-A Action Space Assumption
@@
Action can be categorized into two types: discrete actions, such as controlling grasping and releasing using discrete values [12], and continuous actions, such as the motion of the robot. Discrete actions, by their nature, cannot form a continuous trajectory and therefore do not require trajectory optimization. To address this distinction, we introduce a selection matrix $S\in\mathbb{R}^{D_{c}\times D_{a}}$ that identifies the action dimensions to be considered for trajectory optimization. Specifically, $S$ is used to select: 1) degrees of freedoms (DOFs) that are inherently continuous and capable of forming a trajectory, and 2) DOFs that require trajectory optimization in practical applications. The resulting filtered action variable is expressed as $c=Sa\in\mathbb{R}^{D_{c}}$. In this paper, we filter out the grasping action using $S$, as grasping typically does not require trajectory motion constraints and optimization.
@@
Assumption 1.
Another widely adopted policy action space is $[v_{x},\,v_{y},\,v_{z},\,\dot{r}_{x},\,\dot{r}_{y},\,\dot{r}_{z}]^{\mathrm{T}}\in\mathbb{R}^{6},$ where $\mathbf{v}_{xyz}=[v_{x},v_{y},v_{z}]^{\mathrm{T}}$ denotes the translational velocity, and $\dot{\mathbf{r}}=[\dot{r}_{x},\dot{r}_{y},\dot{r}_{z}]^{\mathrm{T}}$ is the time derivative of the axis–angle rotation vector [12, 26, 36, 32]. Taking the time derivative gives $[\dot{v}_{x},\,\dot{v}_{y},\,\dot{v}_{z},\,\ddot{r}_{x},\,\ddot{r}_{y},\,\ddot{r}_{z}]^{\!\mathrm{T}}.$ The first three elements are explicit linear accelerations, whereas the last three are axis-angle second derivatives, not the physical angular acceleration $\dot{\bm{\omega}}$. Therefore, this type of action space is not covered by Assumption 1. However, we will show that under this type of action representation, constraints on the action derivative can still effectively regulate the robot’s physical motion.
The relationship between the axis-angle representation and the angular acceleration is governed by a Jacobian matrix $\mathbf{J}(\mathbf{r})$, which maps $\dot{\mathbf{r}}$ to the angular velocity $\bm{\omega}$ such that $\bm{\omega}=\mathbf{J}(\mathbf{r})\dot{\mathbf{r}}$. Taking the time derivative, the angular acceleration is given by $\dot{\bm{\omega}}=\mathbf{J}(\mathbf{r})\ddot{\mathbf{r}}+\dot{\mathbf{J}}(\mathbf{r})\dot{\mathbf{r}}$, where $\ddot{\mathbf{r}}$ is the second derivative of the axis-angle vector and $\dot{\mathbf{J}}(\mathbf{r})$ is the time derivative of the Jacobian.
Although $\ddot{\mathbf{r}}$ does not directly represent $\dot{\bm{\omega}}$, constraining $\ddot{\mathbf{r}}$ still effectively regulates $\dot{\bm{\omega}}$. In the ideal case of small rotation angles and slow rotational motions, the term $\dot{\mathbf{J}}(\mathbf{r})\dot{\mathbf{r}}$ becomes negligible, and $\mathbf{J}(\mathbf{r})$ approximates the identity matrix. This yields $\dot{\bm{\omega}}\approx\ddot{\mathbf{r}}$, making constraints on $\ddot{\mathbf{r}}$ nearly equivalent to constraining the true angular acceleration.
In practice, while $\mathbf{J}(\mathbf{r})$ may deviate from the identity matrix, it typically remains bounded. This boundedness is evidenced by the widespread empirical success of using the axis-angle velocity as the rotational action in many imitation learning frameworks, such as those in [12, 26, 36]. These works employ end-to-end learning without any explicit constraints on $\dot{\mathbf{r}}$ and $\ddot{\mathbf{r}}$, yet still achieve rotationally stable behaviors. This empirical stability indicates that the mapping from axis-angle velocity to true angular velocity is well-behaved, further supporting the validity of this action space.
Building on this insight, DiffOG explicitly constrains and smooths $\ddot{\mathbf{r}}$ (Section III-C), enabling finer control over rotational motion and resulting in smoother, more physically consistent behaviors. Therefore, even though this action representation does not strictly satisfy Assumption 1, it still enables effective regularization of angular acceleration. Experimental results in Section V-A confirm that DiffOG led to improved motion for this type of action space.
@@
III-B Overview
In imitation learning, the policy learns the demonstrated task skills through supervised learning on the demonstration dataset, as shown in Fig. 1. The successful execution of a task depends on the policy generating action data that matches the distribution of the demonstration dataset. The objective of supervised learning in this context is to enable the policy to generate action that follows the same distribution as the demonstration dataset. However, traditional policy learning neglects trajectory constraints and smoothness. As a result, post-processing the policy’s output to enforce these properties tends to shift the actions away from the demonstrated distribution, thereby compromising policy performance.
To overcome this challenge, DiffOG leverages the differentiability of the optimization problem to directly learn a trajectory optimizer within a supervised learning framework. This trajectory optimizer is designed to balance objectives including constraint satisfaction, smoothness, and fidelity to the dataset, such that the resulting actions remain within the demonstrated distribution. Moreover, by incorporating transformer into the differentiable optimization layer, DiffOG significantly enhances both the generalization ability and representational power of trajectory optimization, making it applicable to a wide range of tasks, including long-horizon manipulation tasks with high-dimensional action spaces.
@@ III-C Differentiable Trajectory Optimization
In the optimization problem (6), $\mathbf{Q}$ is a weight matrix, $\alpha$ is a scalar, and $d_{\text{max}},d_{\text{min}}$ are constraint bounds for the derivatives of the robot actions. Intuitively, $\alpha$ reflects the degree to which the trajectory smoothness is emphasized. When the sum of the derivatives of the robot actions $\hat{\mathbf{c}}_{t}$ in the generated trajectory $\mathbf{y}_{t}$ is large, it indicates that the trajectory is less smooth. For instance, excessive acceleration or velocity contributes to an increase in the cost function. Therefore, we call $\alpha$ as smoothing weight. Optimization problem (6) aims to solve for an action trajectory that satisfies the constraints on action derivatives while achieving greater smoothness, making it a trajectory optimization problem. $\mathbf{Q}$ also has a intuitive theoretical interpretation, which will be elaborated in detail in Section III-D. In Section III-D, we will delve deeper into the trajectory optimization objectives by incorporating learning-based perspectives.
The optimization problem (6) is a quadratic program, which is a class of optimization problems. In DiffOG, we instantiate this class by proposing a specific, learnable trajectory optimization formulation tailored for imitation learning. Thus far, we have introduced the overall formulation of trajectory optimization in DiffOG. To further enable learning within this optimization layer, it is essential to emphasize a key property: differentiability. A differentiable trajectory optimization module can be seamlessly integrated with neural networks for end-to-end training, enabling the optimization process to acquire stronger representational capacity through data-driven learning. In what follows, we delve into the details of differentiability, beginning with the following proposition:
@@
Because rank($S$) = $D_{c}$, rank($\mathbf{S}$) = $T_{p}D_{c}$. Because $c$ is taken from a subset of $a$, it follows that $D_{a}\geq D_{c}$. Therefore, $\mathbf{S}$ has full rank. Since $\mathbf{Q}$ is symmetric positive definite and the smoothing weight $\alpha\geq 0$, the matrix $\mathbf{Q}+\alpha\mathbf{S}^{\mathrm{T}}\mathbf{A}_{\text{diff}}^{\mathrm{T}}\mathbf{A}_{\text{diff}}\mathbf{S}$ is also symmetric positive definite. The condition $d_{\text{min}}<d_{\text{max}}$ ensures the feasibility of constraints, guaranteeing constraints are not contradictory.
@@ III-D The Learning Objectives of DiffOG
1) The first objective is to make the output $\mathbf{y}_{t}^{\star}$ as close as possible to $\mathbf{a}_{t}$. From the perspective of imitation learning, this makes that $\mathbf{y}_{t}^{\star}$ remain fidelity to demonstration dataset, thereby enabling the actions generated by DiffOG to successfully perform the demonstrated task.
@@ III-E Parameterize $\mathbf{Q}$ with Transformer
Thanks to the well-posed properties, namely feasibility and differentiability, of our proposed trajectory optimization formulation, we are able to incorporate a transformer module to significantly enhance its capacity through a series of careful design choices. This enhancement enables the optimization process to generalize to a wider range of more complex tasks, while preserving interpretability (see Remark 3) and maintaining feasibility and training stability (see Remark 1). We use a transformer encoder $E_{\theta}$ to construct a symmetric positive definite $\mathbf{Q}$, as detailed in Algorithm 1 and Fig. LABEL:fig:firstPage. The encoder used in Algorithm 1 which computes the embedding $\mathbf{e}_{t}=E_{\theta}(\mathbf{a}_{t})$ is shown in Fig. 2. In Algorithm 1, steps 3)–5) construct a lower triangular matrix with positive diagonal entries from the embedding output by the transformer. Additionally, the use of clamping prevents value explosion caused by the exponential operation. Based on Cholesky factorization, Step 6) results in a symmetric positive definite $\mathbf{Q}$. Step 7) further reinforces the symmetry of $\mathbf{Q}$ to mitigate the impact of minor numerical errors introduced during computation, ensuring the desired properties of $\mathbf{Q}$ are preserved. In practice, we clamp the values to the range [-10, 10]. For the small constant $\epsilon$, we use $1\times 10^{-4}$.
@@ III-F Inference
The training of DiffOG is on discrete trajectory chunks. Therefore, to fully constrain the generated trajectory during inference, the forward pass of optimization problem (6) need to add a new constraint.
@@
IV Baselines
where $\mathbf{S}$ is for selecting the the action dimensions to be considered for trajectory optimization. Similarly, we define an output variable
IV-A Greedy Constraint Clipping
The optimized action at time $t+k$ is then updated as
$\hat{c}_{t+k}^{\star}=\hat{c}_{t+k-1}^{\star}+\Delta c_{t+k}^{\text{clip}}.$
IV-B Penalty-Based Trajectory Optimization
$\nabla=\mathbf{v}_{\text{max}}-\mathbf{v}_{\text{min}}.$
${\mathbf{\hat{c}}}_{t}\leftarrow{\mathbf{\hat{c}}}_{t}-\eta\nabla,$
V Experiments
@@
For the hyperparameters of Diffusion Policy and DP3, we use the default parameters provided in their respective works [5, 34]. DiffOG, penalty-based optimization, and constraint clipping use same constraints and same base policy, either DP3 or Diffusion Policy, depending on the type of camera input used for the task. For details on the parameters of DiffOG and trajectory processing baselines, as well as information on the datasets, please refer to Appendix.
@@
Comparison with a Constrained Visuomotor Policy: We present comparative experiments between DiffOG and an constrained visuomotor policy Leto [32], showing that DiffOG achieve superior performance.
@@
Figure 3: We validated DiffOG on 13 different tasks. The 13 tasks span three types, corresponding to the three action spaces discussed in Section III-A.
@@
Base Policy: Diffusion Policy ${1.00}\pm{0.00}$ ${0.98}\pm{0.01}$ ${0.91}\pm{0.01}$ $0.83\pm 0.03$ $0.91\pm 0.01$ $0.84\pm 0.03$
DiffOG Dataset (Ours) $\textbf{1.00}\pm\textbf{0.00}$ $\textbf{0.98}\pm\textbf{0.01}$ $0.87\pm 0.02$ $\textbf{0.82}\pm\textbf{0.02}$ $0.89\pm 0.02$ $\textbf{0.83}\pm\textbf{0.02}$
DiffOG Refine (Ours) $\textbf{1.00}\pm\textbf{0.00}$ $\textbf{0.98}\pm\textbf{0.01}$ $\textbf{0.90}\pm\textbf{0.01}$ $0.81\pm 0.02$ $\textbf{0.91}\pm\textbf{0.03}$ $0.80\pm 0.02$
Constraint Clipping $\textbf{1.00}\pm\textbf{0.00}$ $0.93\pm 0.04$ $0.78\pm 0.03$ $0.59\pm 0.03$ $0.82\pm 0.02$ $0.82\pm 0.02$
Penalty-Based Optimization $\textbf{1.00}\pm\textbf{0.00}$ $0.96\pm 0.02$ $0.76\pm 0.03$ $0.79\pm 0.01$ $0.81\pm 0.02$ $0.81\pm 0.04$
@@
Base Policy: DP3 $0.98\pm 0.01$ $0.77\pm 0.04$ $0.87\pm 0.02$ $1.00\pm 0.00$ $0.70\pm 0.02$
DiffOG Dataset (Ours) $\textbf{0.98}\pm\textbf{0.01}$ $\textbf{0.73}\pm\textbf{0.01}$ $0.89\pm 0.05$ $\textbf{1.00}\pm\textbf{0.00}$ $\textbf{0.70}\pm\textbf{0.03}$
DiffOG Refine (Ours) $\textbf{0.98}\pm\textbf{0.01}$ $0.72\pm 0.02$ $\textbf{0.90}\pm\textbf{0.03}$ $\textbf{1.00}\pm\textbf{0.00}$ $0.68\pm 0.02$
Constraint Clipping $\textbf{0.98}\pm\textbf{0.01}$ $0.70\pm 0.01$ $0.86\pm 0.05$ $\textbf{1.00}\pm\textbf{0.00}$ $0.69\pm 0.02$
Penalty-Based Optimization $\textbf{0.98}\pm\textbf{0.01}$ $0.67\pm 0.03$ $0.86\pm 0.06$ $\textbf{1.00}\pm\textbf{0.00}$ $\textbf{0.70}\pm\textbf{0.03}$
@@
Figure 7: Comparison of trajectory metrics across methods on Robomimic and Meta-World benchmarks. These results are derived from 1500 rollouts (across 50 environments, 3 seeds, and the last 10 checkpoints per seed), providing strong statistical significance. For Robomimic tasks, the larger maximum acceleration and greater standard deviation of acceleration over time indicate less smooth trajectories. Similarly, for Meta-World tasks, this observation applies to the velocity metrics. Since the actions provided by the Robomimic dataset are already normalized [12], the acceleration metrics shown here for Robomimic are also normalized. For details on the calculation of acceleration and velocity metrics, as well as the details to unnormalize the Robomimic metrics, please refer to Appendix.
@@
V-A Benchmark with Baselines Across Tasks
In this section, we demonstrate how DiffOG optimizes the trajectories generated by base policies across 13 tasks. The 13 tasks span three types, corresponding to the three action spaces described in Section III-A. For the end-effector 6-DOF velocity action space, we include five tasks: lift, can, square, tool hang, and transport, all sourced from Robomimic benchmark [12]. For the end-effector linear position action space, we include six tasks: push-T from [5], along with pick place wall, shelf place, disassemble, stick push, and stick pull from Meta-World [33]. Notably, we selected all the “very hard” tasks identified in Meta-World as marked in [34]. Lastly, for the joint angle action space, we include two real-world tasks, arrange desk and move the stack, with data collection and policy experiments conducted using ALOHA [35].
For real-world tasks, we randomize the initial positions and orientations of each object while maintaining a consistent placement layout to test the policy’s generalization ability; details are provided in the supplementary video. For the move the stack task, the primary goal of our testing is to evaluate the impact of trajectory smoothness on task performance. Therefore, instead of randomizing the placement of cups and coasters across the entire table for data collection and policy rollout, we placed them on opposite sides of the table. This reduces the number of demonstrations required for generalization.
The purpose of the experiments are to investigate the impact of applying trajectory processing, which enforces specific constraints on action trajectories, on policy performance across a wide range of tasks. Additionally, we aim to evaluate whether this approach effectively makes action trajectories smoother and more constrained. The experimental results are shown in Tables I, II, III and Figs. 7 and 8. By analyzing the experimental results, we can draw the following conclusions.

### v4

@@ Abstract
Imitation learning-based visuomotor policies excel at manipulation tasks but often produce suboptimal action trajectories compared to model-based methods. Directly mapping camera data to actions via neural networks can result in jerky motions and difficulties in meeting critical constraints, compromising safety and robustness in real-world deployment. For tasks that require high robustness or strict adherence to constraints, ensuring trajectory quality is crucial. However, the lack of interpretability in neural networks makes it challenging to generate constraint-compliant actions in a controlled manner. This paper introduces differentiable policy trajectory optimization with generalizability (DiffOG), a learning-based trajectory optimization framework designed to enhance visuomotor policies. By leveraging the proposed differentiable formulation of trajectory optimization with transformer, DiffOG seamlessly integrates policies with a generalizable optimization layer. DiffOG refines action trajectories to be smoother and more constraint-compliant while maintaining alignment with the original demonstration distribution, thus avoiding degradation in policy performance. We evaluated DiffOG across 11 simulated tasks and 2 real-world tasks. The results demonstrate that DiffOG significantly enhances the trajectory quality of visuomotor policies while having minimal impact on policy performance, outperforming trajectory processing baselines such as greedy constraint clipping and penalty-based trajectory optimization. Furthermore, DiffOG achieves superior performance compared to existing constrained visuomotor policy. For more details, please visit the project website: https://zhengtongxu.github.io/diffog-website/.
@@ II-A Supervised Policy Learning
Supervised policy learning entails training a policy using supervised learning techniques on a dataset of pre-collected demonstrations [16]. Recent advancements in machine learning algorithms and data collection frameworks have substantially advanced the capabilities and effectiveness of supervised policy learning in manipulation applications. For example, various data collection frameworks and devices have been proposed in [6, 36, 8, 22, 21, 25], enabling efficient data collection and the training of effective manipulation policies. To better capture the multi-modal action distributions present in demonstration data, prior work have explored various techniques, such as using diffusion models [5], VQ-VAE [10], and energy-based models [7]. Furthermore, various effective action representations for supervised policy learning have been proposed, such as flow by point tracking [31, 28] and equivariant action representation [26]. Moreover, many work have contributed to training more general and scalable policies on large-scale datasets [14, 15, 3, 12].
@@ III Method
III-A Overview
In imitation learning, the policy learns the demonstrated task skills through supervised learning on the demonstration dataset, as shown in Fig. 1. The successful execution of a task depends on the policy generating action data that matches the distribution of the demonstration dataset. The objective of supervised learning in this context is to enable the policy to generate actions that follow the same distribution as the demonstration dataset. However, traditional policy learning neglects trajectory constraints and smoothness. As a result, post-processing the policy’s output to enforce these properties tends to shift the actions away from the demonstrated distribution, thereby compromising policy performance.
To overcome this challenge, DiffOG leverages the differentiability of the optimization problem to directly learn a trajectory optimizer within a supervised learning framework. This trajectory optimizer is designed to balance objectives including constraint satisfaction, smoothness, and fidelity to the dataset, such that the resulting actions remain within the demonstrated distribution. Moreover, by incorporating a transformer into the differentiable optimization layer, DiffOG significantly enhances both the generalization ability and representational power of trajectory optimization, making it applicable to a wide range of tasks, including long-horizon manipulation tasks with high-dimensional action spaces.
III-B Action Space
@@
Actions fall into two categories: discrete actions, such as controlling grasping and releasing using discrete values [13], and continuous actions, such as the motion of the robot. Discrete actions, by their nature, cannot form a continuous trajectory and therefore do not require trajectory optimization. To address this distinction, we introduce a selection matrix $S\in\mathbb{R}^{D_{c}\times D_{a}}$ that identifies the action dimensions to be considered for trajectory optimization. Specifically, $S$ is used to select: 1) degrees of freedoms (DOFs) that are inherently continuous and capable of forming a trajectory, and 2) DOFs that require trajectory optimization in practical applications. The resulting filtered action variable is expressed as $c=Sa\in\mathbb{R}^{D_{c}}$. In this paper, we filter out the grasping action using $S$, as grasping typically does not require trajectory motion constraints and optimization.
@@ III-C Differentiable Trajectory Optimization
In the optimization problem (6), $\mathbf{Q}$ is a weight matrix, $\alpha$ is a scalar, and $d_{\text{max}},d_{\text{min}}$ are constraint bounds for the derivatives of the robot actions. Intuitively, $\alpha$ reflects the degree to which the trajectory smoothness is emphasized. When the sum of the derivatives of the robot actions $\hat{\mathbf{c}}_{t}$ in the generated trajectory $\mathbf{y}_{t}$ is large, it indicates that the trajectory is less smooth. For instance, excessive acceleration or velocity contributes to an increase in the cost function. Therefore, we call $\alpha$ as smoothing weight. Optimization problem (6) aims to solve for an action trajectory that satisfies the constraints on action derivatives while achieving greater smoothness, making it a trajectory optimization problem. $\mathbf{Q}$ also has a intuitive theoretical interpretation, which will be elaborated in detail in Section III-D. In Section III-D, we will dive deeper into the trajectory optimization objectives by incorporating learning-based perspectives.
The optimization problem (6) is a quadratic program. In DiffOG, we instantiate this class by proposing a specific, learnable trajectory optimization formulation tailored for imitation learning. Thus far, we have introduced the overall formulation of trajectory optimization in DiffOG. To further enable learning within this optimization layer, it is essential to emphasize a key property: differentiability. A differentiable trajectory optimization module can be seamlessly integrated with neural networks for end-to-end training, enabling the optimization process to acquire stronger representational capacity through data-driven learning. In what follows, we dive into the details of differentiability, beginning with the following proposition:
@@
[…]
@@ III-D The Learning Objectives of DiffOG
1) The first objective is to make the output $\mathbf{y}_{t}^{\star}$ as close as possible to $\mathbf{a}_{t}$. From the perspective of imitation learning, this keeps $\mathbf{y}_{t}^{\star}$ faithful to the demonstrations, thereby enabling the actions generated by DiffOG to successfully perform the demonstrated task.
@@ III-E Parameterize $\mathbf{Q}$ with Transformer
Thanks to the well-posed properties, namely feasibility and differentiability, of our proposed trajectory optimization formulation, we are able to incorporate a transformer module to significantly enhance its capacity through a series of careful design choices. This enhancement enables the optimization process to generalize to a wider range of more complex tasks, while preserving interpretability (see Remark 3) and maintaining feasibility and training stability (see Remark 1). We use a transformer encoder $E_{\theta}$ to construct a symmetric positive definite $\mathbf{Q}$, as detailed in Algorithm 1 and Fig. LABEL:fig:firstPage. The encoder used in Algorithm 1 which computes the embedding $\mathbf{e}_{t}=E_{\theta}(\mathbf{a}_{t})$ is shown in Fig. 2. In Algorithm 1, steps 3)–5) construct a lower triangular matrix with positive diagonal entries from the embedding output by the transformer. Additionally, the use of clamping prevents numerical explosion caused by exponentiation. Based on Cholesky factorization, Step 6) results in a symmetric positive definite $\mathbf{Q}$. Step 7) further reinforces the symmetry of $\mathbf{Q}$ to mitigate the impact of minor numerical errors introduced during computation, ensuring the desired properties of $\mathbf{Q}$ are preserved. In practice, we clamp the values to the range [-10, 10]. For the small constant $\epsilon$, we use $1\times 10^{-4}$.
@@ III-F Inference
The training of DiffOG is on discrete trajectory chunks. Therefore, to fully constrain the generated trajectory during inference, the forward pass of optimization problem (6) needs to add a new constraint.
@@
IV Experiments
@@
For the hyperparameters of Diffusion Policy and DP3, we use the default parameters provided in their respective work [5, 35]. DiffOG, penalty-based optimization, and constraint clipping use same constraints and same base policy, either DP3 or Diffusion Policy, depending on the type of camera input used for the task. For details on the parameters of DiffOG and trajectory processing baselines, as well as information on the datasets, please refer to Appendix.
@@
Comparison with a Constrained Visuomotor Policy: We present comparative experiments between DiffOG and an constrained visuomotor policy Leto [33], showing that DiffOG achieves superior performance.
Inference Time-varying Constraint Experiment: In this experiment, the policy is trained with a fixed constraint bound but evaluated with time-varying bounds during inference, demonstrating DiffOG’s capability to generalize zero-shot to unseen constraint settings.
Residual Policy Baseline with Smoothness Penalty: In this experiment, we train a residual policy by learning an action correction on top of base policy, with an additional smoothness penalty in the loss. By comparing the performance of DiffOG with this baseline, we further highlight how DiffOG’s structure leads to better alignment with task requirements and stronger overall performance, making it a more principled and generalizable method.
@@
Figure 3: We validated DiffOG on 13 different tasks. The 13 tasks span three types of action space.
@@
[…]
@@
Base Policy: DP3 0.98±0.010.98$\pm$ 0.010.98 ± 0.01 0.77±0.040.77$\pm$ 0.040.77 ± 0.04 0.87±0.020.87$\pm$ 0.020.87 ± 0.02 1.00±0.001.00$\pm$ $0$ 0.00 0.70±0.020.70$\pm$ 0.020.70 ± 0.02
[…]
@@
Figure 7: Comparison of trajectory metrics across methods on Robomimic and Meta-World benchmarks. These results are derived from 1500 rollouts (across 50 environments, 3 seeds, and the last 10 checkpoints per seed), providing strong statistical significance. Note that these values do not represent simple variations around a common mean and therefore cannot be interpreted in the same way as conventional mean $\pm$ standard deviation error bars, as they are not computed from a single rollout trajectory. We report them separately to highlight distinct aspects of trajectory quality. For Robomimic tasks, the larger maximum acceleration and greater standard deviation of acceleration over time indicate less smooth trajectories. Similarly, for Meta-World tasks, this observation applies to the velocity metrics. Since the actions provided by the Robomimic dataset are already normalized [13], the acceleration metrics shown here for Robomimic are also normalized. For details on the calculation of acceleration and velocity metrics, as well as the details to unnormalize the Robomimic metrics, please refer to Appendix.
@@
IV-A Benchmark with Baselines Across Tasks
In this section, we demonstrate how DiffOG optimizes the trajectories generated by base policies across 13 tasks. The 13 tasks span three types of action space. See Appendix C for more details.
The purpose of the experiments is to investigate the impact of applying trajectory processing, which enforces specific constraints on action trajectories, on policy performance across a wide range of tasks. Additionally, we aim to evaluate whether this approach effectively makes action trajectories smoother and more constrained. The experimental results are shown in Tables I, II, III and Figs. 7 and 8.
We compute evaluation metrics based on 1500 rollouts (last 10 checkpoints $\times$ 3 seeds $\times$ 50 environment conditions) to ensure statistical significance. This is important for evaluating both task success rate, which indicates whether the method still effectively captures the distribution of demonstration data under constraints, and motion-related characteristics, which reflects trajectory quality. By analyzing the experimental results, we can draw the following conclusions.